# Gauging the Standard Model $1$-form symmetry via gravitational instantons

### v1

@@ A.1 Mathematical preliminaries
We shall adopt the convention
@@ A.2 The Eguchi-Hanson instanton
where $\bm{x}=(x_{1},x_{2},x_{3})$ are Cartesian coordinates, and $\tau$ may be interpreted as a Euclidean time coordinate. The functions $V(\bm{x})$ and $\bm{w}(\bm{x})=(w_{1}(\bm{x}),w_{2}(\bm{x}),w_{3}(\bm{x}))$ can be viewed, respectively, as graviscalar potential and gracivector potential sourced by a distribution of dyons carrying the same electric and magnetic charges $(e,m)=(1,\pm 1)$. Specifically, we take the scalar potential to be
@@
with velibians given by
@@
In particular, at large $r$, flat $\mathbb{R}^{4}$ approaches the 3-sphere $S^{3}$, parametrized by the Euler angles $(\theta,\varphi,\psi)$ over the specified domain. Comparing this to the EH space as $r\rightarrow\infty$, we observe that the angular part of the metric approaches that of a squashed 3-sphere. However, in the EH geometry, the coordinate $\psi$ has half the range it does in flat space, namely $\Delta\psi=2\pi$ instead of $4\pi$. This implies that the asymptotic boundary of the EH space is not $S^{3}$ but rather the quotient $S^{3}/\mathbb{Z}^{2}=\mathbb{RP}^{3}$. We note that the volume $3$-form on $\mathbb{RP}^{3}$ is given by
@@
and we remind the reader that we use a convention where the wedge product $e^{0}\wedge e^{1}\wedge e^{2}\wedge e^{3}$ is positive since $\epsilon_{0123}=1$.
@@ Appendix B Fermions in the EH background: fermion zero modes
Using $\hat{\cal J}_{-}|j,m^{\prime}=-j,m\rangle=\hat{\cal J}_{+}|j,m^{\prime}=j,m\rangle=0$, Eq. (B) simplifies to (we do not ditnguish between $j_{1}$ and $j_{2}$, referring to both total angular momentum by $j$)
@@
We take the path $\ell_{a}$ that connects two points $(r=a,\theta=0,\varphi=0,\psi)$ and $(r=a,\theta=0,\varphi=0,\psi+2\pi)$ on the bolt $\mathbb{S}^{2}$, which is contractible. Using (145, 166, 150), and $f^{-1}(r=a)=0$, yields
$\exp\left[-\frac{1}{2}\int_{\ell_{a}}\omega_{ab}\Sigma^{ab}\right]$ $=$ $\exp\left[-i\int_{\ell_{a}}\left[\begin{array}[]{cc}\frac{\sigma_{x}\hat{\sigma}_{x}}{f}+\frac{\sigma_{y}\hat{\sigma}_{y}}{f}+(2-f^{-2})\sigma_{z}\hat{\sigma}_{z}&0\\ 0&0\end{array}\right]\right]=I_{4}\,$, (194)
$\exp\left[i\int_{\ell_{a}}A\right]$ $=$ $e^{i\pi{\cal C}}\,.$ (195)
Notice that the gravitational holonomy is a trivial phase upon parallel transporting the spinor along $\ell_{a}$. Thus, under gravity and gauge fields, parallel transporting the spinor along the path $\ell_{a}$ gives
$\Psi(r=a,\theta,\varphi,\psi+2\pi)$ $=e^{i\pi{\cal C}}\Psi(r=a,\theta,\varphi,\psi)\,.$ (196)
@@
The fermion zero modes are well-defined on EH space, i.e., they are unambiguously defined on a contactible path, if and only if the r.h.s. of (196) and (201) are identical.
@@
We use the boundary conditions at the bolt (196) to determine the number of globally well-defined zero modes. We distinguish two cases.
Case I: $\cal C$} is even, $C=2p\geq 0$, the gauge holonomy at the bolt gives a trivial phase, and thus, one must restrict the values of $j$ to integers. In this case, the number of normalizable zero modes is
@@
Case II: $\cal C$} is odd, ${\cal C}=2p+1\geq 1$, the gauge holonomy at the bolt gives an $e^{i\pi}$ phase, and thus, one must restrict $j$ to half-integers. The total number of normalizable zero modes in this case is
@@ Appendix C The Atiyah-Patodi-Singer index theorem
Here, $\Theta$ is the second fundamental form computed at the boundary of $\mathbb{M}$ and $\eta$ is the eta-inavriant defined as
@@
The contribution from the $U(1)$ and gravitational backgrounds were already computed above:
@@ Appendix D Fermions in the background of $PSU(2)$ gauge field
$A_{SU(2)}=m_{(2)}H\nu=m_{(2)}\left[\begin{array}[]{cc}\frac{1}{2}&0\\ 0&-\frac{1}{2}\end{array}\right]\frac{\sigma_{z}a^{2}}{r^{2}}\,$, (237)
@@ D.1 Fundamental fermions
However, as in the $U(1)$ case, not all normalizable modes possess a spin structure upon parallel transport along closed paths. Under parallel transport from the point $(r=a,\theta=0,\varphi=0,\psi)$ to $(r=a,\theta=0,\varphi=0,\psi+2\pi)$, the Dirac spinor aquires a $SU(2)$ holonomy:
@@ D.2 Adjoint fermions
also consistent with the calculations of the APS index. In particular, for $m_{(2)}=2$, a background gauge field carrying a tological charge $Q_{(2)}=\frac{1}{2}$, there are $2$ adjoint zero modes.

### v2

@@ A.1 Mathematical preliminaries
We shall adopt the convention (this defines our convention for the frame orientation)
@@ A.2 The Eguchi-Hanson instanton
where $\bm{x}=(x_{1},x_{2},x_{3})$ are Cartesian coordinates, and $\tau$ may be interpreted as a Euclidean time coordinate. The functions $V(\bm{x})$ and $\bm{w}(\bm{x})=(w_{1}(\bm{x}),w_{2}(\bm{x}),w_{3}(\bm{x}))$ can be viewed, respectively, as graviscalar potential and gravivector potential sourced by a distribution of dyons carrying the same electric and magnetic charges $(e,m)=(1,\pm 1)$. Specifically, we take the scalar potential to be
@@
with vielbeins given by
@@
In particular, at large $r$, flat $\mathbb{R}^{4}$ approaches the 3-sphere $S^{3}$, parametrized by the Euler angles $(\theta,\varphi,\psi)$ over the specified domain. Comparing this to the EH space as $r\rightarrow\infty$, we observe that the angular part of the metric approaches that of a 3-sphere. However, in the EH geometry, the coordinate $\psi$ has half the range it does in flat space, namely $\Delta\psi=2\pi$ instead of $4\pi$. This implies that the asymptotic boundary of the EH space is not $S^{3}$ but rather the quotient $S^{3}/\mathbb{Z}^{2}=\mathbb{RP}^{3}$. We note that the volume $3$-form on $\mathbb{RP}^{3}$ is given by
@@
and we use a convention where the wedge product $e^{0}\wedge e^{1}\wedge e^{2}\wedge e^{3}$ is positive since for the orientation frame we chose $\epsilon_{0123}=1$. With the convention that $e^{0}\wedge e^{1}\wedge e^{2}\wedge e^{3}>0$, the positive volume form in the $(r,\theta,\varphi,\psi)$ coordinates is $dr\wedge d\theta\wedge d\varphi\wedge\psi$; thus, the coordinate orientation we choose is opposite to the frame orientation. This choice makes it easier to apply the index theorem without having to carry an overall negative sign.
@@ Appendix B Fermions in the EH background: fermion zero modes
Using $\hat{\cal J}_{-}|j,m^{\prime}=-j,m\rangle=\hat{\cal J}_{+}|j,m^{\prime}=j,m\rangle=0$, Eq. (B) simplifies to (we do not distinguish between $j_{1}$ and $j_{2}$, referring to both total angular momentum by $j$)
@@
We take the path $\ell_{r\approx a}$ that connects two points $(r\approx a,\theta=0,\varphi=0,\psi)$ and $(r\approx a,\theta=0,\varphi=0,\psi+2\pi)$ near the bolt $\mathbb{S}^{2}$, which is contractible. Using (145, 166, 150), and $f^{-1}(r\approx a)=0$, yields
$\exp\left[-\frac{1}{2}\int_{\ell_{r\approx a}}\omega_{ab}\Sigma^{ab}\right]$ $=$ $\exp\left[-i\int_{\ell_{r\approx a}}\left[\begin{array}[]{cc}\frac{\sigma_{x}\hat{\sigma}_{x}}{f}+\frac{\sigma_{y}\hat{\sigma}_{y}}{f}+(2-f^{-2})\sigma_{z}\hat{\sigma}_{z}&0\\ 0&0\end{array}\right]\right]=I_{4}\,$, (194)
$\exp\left[i\int_{\ell_{r\approx a}}A\right]$ $=$ $e^{i\pi{\cal C}}\,.$ (195)
Notice that the gravitational holonomy is a trivial phase upon parallel transporting the spinor along $\ell_{\approx a}$. Thus, under gravity and gauge fields, parallel transporting the spinor along the path $\ell_{\approx a}$ gives
$\Psi(r\approx a,\theta,\varphi,\psi+2\pi)$ $=e^{i\pi{\cal C}}\Psi(r\approx a,\theta,\varphi,\psi)\,.$ (196)
@@
The fermion zero modes are well-defined on EH space, i.e., they are unambiguously defined on a contractible path, if and only if the r.h.s. of (196) and (201) are identical.
@@
[…]
In conclusion, we use the boundary conditions near the bolt (196) to determine the number of globally well-defined zero modes. We distinguish two cases.
Case I: $\cal C$} is even, $C=2p\geq 0$, the gauge holonomy near the bolt gives a trivial phase, and thus, one must restrict the values of $j$ to integers. In this case, the number of normalizable zero modes is
@@
Case II: $\cal C$} is odd, ${\cal C}=2p+1\geq 1$, the gauge holonomy near the bolt gives an $e^{i\pi}$ phase, and thus, one must restrict $j$ to half-integers. The total number of normalizable zero modes in this case is
@@ Appendix C The Atiyah-Patodi-Singer index theorem
Here, $\Theta$ is the second fundamental form computed at the boundary of $\mathbb{M}$ and $\eta$ is the $\eta$-invariant defined as
@@
Although the EH space is noncompact, its asymptotic region has the structure of a cylindrical end $\mathbb{R}_{+}\times\mathbb{RP}^{3}$. To apply the APS index theorem in this ALE background, we introduce a large radial cutoff $r=R$ (a regulator) and restrict to the compact manifold $\mathbb{M}_{R}$ with boundary $\partial\mathbb{M}_{R}\simeq\mathbb{RP}^{3}$ at $r=R$. On $\partial\mathbb{M}_{R}$, we keep fixed the induced metric and the pullback of the background gauge connection, including the $\mathbb{Z}_{2}$ holonomy information, and we choose the compatible (twisted) spin structure. The APS boundary condition, which is a non-local condition, restricts the boundary value of the spinor to lie in the subspace spanned by eigenmodes of the three-dimensional Dirac operator on $\mathbb{RP}^{3}$ (with flat bundle) whose eigenvalues have a prescribed sign, and ensures that the Dirac operator is self-adjoint at the boundary. The index computed from (C) on $\mathbb{M}_{R}$ is independent of $R$ (since it is a topological invariant independent of the regulator), so we can take the limit $R\to\infty$ and regard the result as the index of the Dirac operator on the full EH space with an asymptotically cylindrical end. Notice that the use of the APS non-local boundary conditions is totally acceptable when it comes to counting the zero modes, as it only provides a convenient mathematical device to guarantee that the Dirac operator is self-adjoint. For a physicist’s friendly review of this procedure on a general manifold with boundaries, we refer the reader to Ninomiya:1984ge .
The contribution from the $U(1)$ and gravitational backgrounds to the APS index were already computed above:
@@ Appendix D Fermions in the background of $PSU(2)$ gauge field
$A_{SU(2)}=m_{(2)}H\nu\frac{\sigma_{z}a^{2}}{r^{2}}=m_{(2)}\left[\begin{array}[]{cc}\frac{1}{2}&0\\ 0&-\frac{1}{2}\end{array}\right]\frac{\sigma_{z}a^{2}}{r^{2}}\,$, (237)
@@ D.1 Fundamental fermions
However, as in the $U(1)$ case, not all normalizable modes possess a spin structure upon parallel transport along closed paths. Under parallel transport from the point $(r\approx a,\theta=0,\varphi=0,\psi)$ to $(r\approx a,\theta=0,\varphi=0,\psi+2\pi)$, the Dirac spinor acquires a $SU(2)$ holonomy:
@@ D.2 Adjoint fermions
also consistent with the calculations of the APS index. In particular, for $m_{(2)}=2$, a background gauge field carrying a topological charge $Q_{(2)}=\frac{1}{2}$, there are $2$ adjoint zero modes.